# Metaverse Identity: Core Principles and Critical Challenges

### v1

@@ 4.3 Privacy and Identity Management Challenges
The metaverse presents unprecedented challenges to user privacy. A study by Nair et al., 2023a , involving over 50,000 players of the popular VR Game ’Beat Saber’, highlights these concerns. Analysis of 2.5 million VR motion dataset through machine learning algorithms reveals that merely 100 seconds of data can uniquely identify a user with over 94.33% accuracy, and remarkably, only 2 seconds are needed to identify half of the users (Nair et al., 2023a, ). Further research indicates that this motion data can accurately infer a broad spectrum of personal characteristics, including bio-metric information such as height and wingspan, as well as demographic details like age and gender. It can even predict the user’s country of origin and clothing type. Most concerning is the accurate capability to discern the presence of mental and physical disabilities (Nair et al., 2023b, ). This level of precision in identification intensifies when motion data are coupled with other tracked data within the metaverse. This situation suggest that anonymity in the metaverse may be impossible (Rosenberg,, 2023).
@@
Contrary to traditional privacy measures that do not necessitate sharing sensitive data like fingerprints, the metaverse inherently involves the dissemination of motion data–a fundamental aspect of interaction that must be shared in real-time with all participants. Furthermore, unique motion patterns analyzed across various contexts–be it professional environments, social interactions, or private setting–facilitate the straightforward identification of individuals. Additionally, the inherent link between our physical movements and our motion data allows machine learning algorithms to correlate VR actions with real-world surveillance footage, leading to potential real-world tracking and identification (Nair et al.,, 2024).
@@
In summary, without the development of innovative protective measures, maintaining privacy in the metaverse could be an insurmountable challenge, presenting significant and potentially transformative implications for identity management. To effectively address the unique challenges the metaverse poses to privacy, targeted measures are necessary both technologically and at the policy level. From a technological standpoint, various sophisticated methods, including local epsilon-differential privacy, adversarial machine learning, Trusted Execution Environments (TEE), and Secure Multi-Party Computation (MPC), can enhance user privacy (Nair et al.,, 2024). However, these technologies can sometimes compromise user experience or remain underdeveloped (Nair et al.,, 2024). No comprehensive and fully mature technical solutions currently exist, underscoring the importance of establishing robust privacy regulations to complement and guide the development while fostering the development of privacy-enhancing technologies. Addressing the metaverse’s unique privacy challenges requires a dual approach: continuously improving privacy-enhancing technologies to keep pace with the metaverse’s rapid evolution, and developing targeted privacy protection measures tailored to its immersive, three-dimensional nature. For example, governing privacy in three-dimensional spaces necessitates distinct data processing and privacy protection measures for public areas (akin to offices) and private spaces (akin to homes). Such differentiation is crucial to ensure that privacy regulations in the metaverse reflect varying real-world privacy expectations, adhering to the principle of Equivalence and Alignment. By adopting this nuanced approach to privacy governance, we can foster a more secure and trustworthy metaverse that respects users’ privacy rights while promoting innovation and growth.users.
@@ 4.4 Challenges of Deepfake and Synthetic Identities
One development trend in the metaverse is to enhance user experiences by making them more realistic, integrated, and immersive (Gonçalves et al.,, 2022), illustrating the principle of Fusion. This trend manifests in several key areas: the application of advanced algorithms such as Generative Adversarial Network(GANs) has significantly enhanced the realism in rendering virtual avatars and digital entities (Tewari et al.,, 2020; Saxena and Cao,, 2021; Teotia et al.,, 2024; Koh et al.,, 2024). Innovative methods by researchers in facial expression (Gonzalez-Franco et al.,, 2020), motion capture (Jung et al.,, 2022), and behavioral modeling (Ahuja et al.,, 2021) have similarly advanced the interactivity and authenticity of avatars and digital entities (Wen et al.,, 2021; Fysh et al.,, 2022). Moreover, there have been considerable advances in the generating and rendering of virtual environments (Wang et al.,, 2021; Fink et al.,, 2023). The rapid evolution of AI, particularly in the field of generative AI, has simplified the creation of highly realistic, indistinguishable digital content, thus enhancing the immersive experience of the metaverse (Ramalingam et al.,, 2023; Murala and Panda,, 2023; Qayyum et al.,, 2024).
@@
However, this technological advancement also poses significant security risks, such as the potential for creating deepfakes or synthesizing untraceable identities, thereby raising serious security concerns (Tariq et al.,, 2023; Wu et al.,, 2023). Deepfakes, manipulated media created using AI and deep learning designed to deceive viewers, have become increasingly prevalent, sparking significant societal discussion (Mirsky and Lee,, 2021). For instance, in 2023, an AI-generated image of the Pentagon explosion went viral, temporarily impacting the U.S. stock market (Alba,, 2023). Generative AI is transforming the creation and expression of digital media content, while the metaverse is poised to redefine the mechanism of content distribution of content distribution, experience, and interaction (Alba,, 2023; WEF, 2023b, ). It is anticipated that the integration of AI with VR and AR technologies will magnify their impacts within the metaverse (WEF, 2023b, ). These technologies facilitate the creation of hyper-realistic simulations, the development of digital entities and scenes, and enable deeper, more emotionally resonant interactions between users or with digital entities.
@@
Consequently, this technological expansion broadens the scope of threats like deepfakes (Tariq et al.,, 2023; Wu et al.,, 2023). In the absence of stringent regulatory oversight within the metaverse, users might exploit personalized identity markers-such as those derived from actual facial features and body forms-to commit identity theft or create synthetic identities. This poses significant risks of severe security and privacy violations, potentially leading to more profound forms of deception, defamation, and threats within the metaverse. For instance, the application of deepfake technology in virtual workplaces to fabricate digital clones of supervisors or colleagues could undermine internal trust, induce confusion by issuing deceptive work directives, or facilitate the dissemination of misinformation or propaganda, thereby impacting critical decisions (Tariq et al.,, 2023). Regrettably, existing deepfake detection methodologies are primarily tailored to address content in the physical world, such as images and videos, and do not yet adequately account for the unique challenges presented by the metaverse (Tariq et al.,, 2023). According to Wu et al., (2023), there is a pressing need for deepfake recognition and prevention mechanisms specifically designed for the metaverse. Looking forward, addressing these challenges will necessitate a dual approach: the development of robust technological solutions for identity verification and content authenticity, and the establishment of clear regulatory frameworks that delineate legal boundaries for virtual avatars/digital content and address critical issues such as privacy protection and misuse prevention. Only through a coordinated effort can the metaverse be ensured to be a relatively safe and trustworthy virtual space.
@@ 4.5 Identity Fragmentation and Psychological Well-being Challenges
The principle of Expansiveness elucidates that the metaverse offers unparalleled opportunities for diverse and expansive identity representation, enabling individuals to transcend the limitations associated with physical world. While this diversity in representation can enhance creativity and inclusivity, as exemplified by the Avatar Robot Cafe (Hatada et al.,, 2024) case where marginalized groups gained equal participatory opportunities, it also poses significant risks such as identity fragmentation and confusion. Recent research by Yang et al., 2024b , involving a two-wave panel study on the VRChat platform, corroborates these concerns, indicating that while expansion of identity representation can boost self-esteem and life satisfaction, it may compromise these benefits when it leads to inconsistencies in self-concept. A poignant historical example is the 2000 incident where a child, after prolonged engagement with a video game, began to experience severe identity dissonance. Believing himself to be the game’s protagonist, he ventured out at night to "fight enemies and save the princess," mimicking the game’s narrative. This blurring of reality and fantasy necessitated medical intervention, to help reorient his sense of identity and reality (Banerji,, 2023).
@@
As the metaverse progresses, addressing these psychological impacts and implementing protective measures for identity integrity, particularly among children and vulnerable populations, becomes imperative. Identity fragmentation within the metaverse can significantly impair users’ mental health and social relationships. The ability to easily create and switch between diverse, unintegrated identities may reduce reliance on real-world personas, exacerbating identity fragmentation issues (Yang et al., 2024b, ). These ’consequence free’ virtual spaces (Turkle,, 1999), which may diverge sharply from reality, pose specific risks. Such environments can lead to self-disalignment, where users perceive their virtual identities as superior, increasing their engagement with virtual communities and causing alienation from the physical world (Banerji,, 2023). This alienation can compel individuals, particularly children and adolescents, to increasingly immerse themselves in the metaverse, potentially leading to social withdrawal, depression, and antisocial behavior (Koohsari et al.,, 2023). Considering the known adverse effects of social media, video games, and mobile devices, the immersive and escapist nature of the metaverse requires careful scrutiny to prevent its potentially detrimental impacts on psychological and social well-being (Banerji,, 2023; Koohsari et al.,, 2023).
@@
Research on multiple cultural identities has shown that the mere presence of multiple identity representations is not inherently problematic; rather, the critical factor is relationships, integration and management of these identity representations (Yang et al., 2024b, ). As the metaverse evolves and identity-related issues become more pronounced, further research and practical interventions are essential. Future studies should focus on assisting users in managing their multiple identity representations and preventing identity fragmentation to mitigate associated psychological and social challenges. Given that these concepts are still in their nascent stages, proactively recognizing and preparing for identity fragmentation is crucial for addressing the emerging challenges of the metaverse effectively.
@@ 5 Conclusion and Limitation
The metaverse, an emerging realm that merges virtual and real elements, significantly influences identity expression, social interaction, and human development trajectories. Central to the metaverse is the construction of identity, which involves representation, data management and identification of avatars and other digital entities (WEF,, 2024). These elements are intimately connected to users’ sense of belonging, privacy, security, and trust. Despite its significance, scholarly research on metaverse identities is still nascent and lacks a comprehensive theoretical framework. This paper conducts a critical review of the existing literature to explore the Metaverse Identity concept from both multidisciplinary and perspective viewpoints, identifying the research gap. Unlike traditional internet settings, the metaverse offers more immersive and varied experiences, transforming identity from a mere theoretical construct into a tangible, experiential reality. This shift accentuates the tension between self-expression and privacy protection.
@@
Metaverse identities are characterized by their multilayered, dynamically evolving, and diverse nature, which grants users unprecedented freedom in self-expression but also poses substantial privacy and security risks. Addressing these complexities, this paper proposes two core principles: Equivalence and Alignment and Fusion and Expansiveness. The first principle argues for consistency between metaverse and real-world identities in terms of behavioral norms and social standards, which is essential for developing conduct guidelines and protecting rights. The second principle stresses the need for deep integration and extensive expansion of metaverse identities, breaking beyond real-world constraits to meet diverse needs and promote inclusive participation. Effective governance in the metaverse necessitates a dynamic balance between these principles, ensuring fairness while encouraging diverse expressions and innovative developments. Our analysis identifies five key challenges to the deveolopment of metaverse identities: interoperability, legal boundaries, privacy and identity management issues, and the risks associated with deep fakes & synthetic identities, and identity fragmentation impacting psychological health. Building on the principles outlined, this paper offers strategic recommendations to address these challenges.
@@
However, this study exhibits several significant limitations. Firstly, the representativeness of the case studies needs enhancement. This paper discusses instances such as virtual sexual assault and Avatar Robot Cafe, yet the case selection is narrow and largely confined to specific scenarios. Given the metaverse’s diverse and complex ecosystem, our analysis might not fully capture the spectrum of identity practices. Future research should broaden the scope of case studies to improve the findings’ generalizability and relevance. Secondly, the empirical foundation of this study requires strengthening. Metaverse identity research, still in its infancy, largely lacks comprehensive data on user behavior and psychological responses. Currently supported primarily by literature review and qualitative analysis, this study highlights the need for more rigorous empirical research. Future efforts should include observational, survey, and experimental methodologies to explore identity construction behaviors and perceptions in the metaverse more deeply, thereby solidifying the theoretical framework. Lastly, the theoretical principles introduced—Equivalence and Alignment, and Fusion and Expansiveness—need further refinement and operationalization. While these principles offer valuable insights for regulating identity in the metaverse, they are proposed from a broad perspective and remain abstract. The practical application and policy-making processes might encounter challenges in operationalizing these principles. Future research should refine these two principles and develop more detailed, actionable guidelines to inform the governance and design of metaverse identity framework.
@@
Looking forward, the metaverse, as an emerging socio-technical ecosystem in the digital age (Lee et al.,, 2021; Wang et al.,, 2022; Lee et al., 2023b, ), is poised to significantly transform human life and developmental practices. The manner in which individuals construct and manage their identities between the physical world and metaverse is critical for sustainable human development. Addressing this issue requires a unified approach from academia, industry, and policy sectors to promote sustained interdisciplinary and multi-perspective research, aimed at fostering a robust and inclusive metaverse environment. Such collaborative initiatives are essential to provide both intellectual and practical contributions to the evolution of this new era.

### v2

@@ 4.3 Privacy and Identity Management Challenges
The metaverse presents unprecedented challenges to user privacy. A study by  Nair et al., 2023a , involving over 50,000 players of the popular VR Game ‘Beat Saber’, highlights these concerns. Analysis of 2.5 million VR motion datasets through machine learning algorithms reveals that merely 100 seconds of data can uniquely identify a user with over 94.33% accuracy (Nair et al., 2023a, ). Further research indicates that this motion data can accurately infer a broad spectrum of personal characteristics, including bio-metric information such as height and wingspan, as well as demographic details like age and gender. It can even predict the user’s country of origin and clothing type. Most concerning is the accurate capability to discern the presence of mental and physical disabilities (Nair et al., 2023b, ). This level of precision in identification intensifies when motion data are coupled with other tracked data within the metaverse. This situation suggest that anonymity in the metaverse may be impossible (Rosenberg,, 2023).
@@
Contrary to traditional privacy measures that do not necessitate sharing sensitive data like fingerprints, the metaverse inherently involves the dissemination of motion data – a fundamental aspect of interaction that must be shared in real-time with all participants. Furthermore, unique motion patterns analyzed across various contexts – whether professional environments, social interactions, or private settings – facilitate the straightforward identification of individuals. Additionally, the intrinsic link between our physical movements and our motion data allows machine learning algorithms to correlate VR actions with real-world surveillance footage, leading to potential real-world tracking and identification (Nair et al.,, 2024).
@@
In summary, without the development of innovative protective measures, maintaining privacy in the metaverse could be an insurmountable challenge, presenting significant and potentially transformative implications for identity management. To effectively address the unique challenges the metaverse poses to privacy, targeted measures are necessary both technologically and at the policy level. From a technological standpoint, various sophisticated methods, including local epsilon-differential privacy, adversarial machine learning, Trusted Execution Environments (TEE), and Secure Multi-Party Computation (MPC), can enhance user privacy (Nair et al.,, 2024). However, these technologies can sometimes compromise user experience or remain underdeveloped (Nair et al.,, 2024). No comprehensive and fully mature technical solutions currently exist, underscoring the importance of establishing robust privacy regulations to complement and guide the development while fostering the development of privacy-enhancing technologies. Addressing the metaverse’s unique privacy challenges requires a dual approach: continuously improving privacy-enhancing technologies to keep pace with the metaverse’s rapid evolution, and developing targeted privacy protection measures tailored to its immersive, three-dimensional nature. For example, governing privacy in three-dimensional spaces necessitates distinct data processing and privacy protection measures for public areas (akin to offices) and private spaces (akin to homes). Such differentiation is crucial to ensure that privacy regulations in the metaverse reflect varying real-world privacy expectations, adhering to the principle of Equivalence and Alignment. By adopting this nuanced approach to privacy governance, we can foster a more secure and trustworthy metaverse that respects users’ privacy rights while promoting innovation and growth.
@@ 4.4 Challenges of Deepfake and Synthetic Identities
One development trend in the metaverse is to enhance user experiences by making them more realistic, integrated, and immersive (Gonçalves et al.,, 2022), illustrating the principle of Fusion. This trend manifests in several key areas: the application of advanced algorithms such as Generative Adversarial Networks(GANs) has significantly enhanced the realism in rendering virtual avatars and digital entities (Tewari et al.,, 2020; Saxena and Cao,, 2021; Teotia et al.,, 2024; Koh et al.,, 2024). Innovative methods by researchers in facial expression (Gonzalez-Franco et al.,, 2020), motion capture (Jung et al.,, 2022), and behavioral modeling (Ahuja et al.,, 2021) have similarly advanced the interactivity and authenticity of avatars and digital entities (Wen et al.,, 2021; Fysh et al.,, 2022). Moreover, there have been considerable advances in the generating and rendering of virtual environments (Wang et al.,, 2021; Fink et al.,, 2023). The rapid evolution of AI, particularly in the field of generative AI, has simplified the creation of highly realistic, indistinguishable digital content, thus enhancing the immersive experience of the metaverse (Ramalingam et al.,, 2023; Murala and Panda,, 2023; Qayyum et al.,, 2024).
@@
However, this technological advancement also poses significant security risks, such as the potential for creating deepfakes or synthesizing untraceable identities, thereby raising serious security concerns (Tariq et al.,, 2023; Wu et al.,, 2023). Deepfakes, manipulated media created using AI and deep learning designed to deceive viewers, have become increasingly prevalent, sparking significant societal discussion (Mirsky and Lee,, 2021). For instance, in 2023, an AI-generated image of the Pentagon explosion went viral, temporarily impacting the U.S. stock market (Alba,, 2023). Generative AI is transforming the creation and expression of digital media content, while the metaverse is poised to redefine the mechanism of content distribution, experience, and interaction (Alba,, 2023; WEF, 2023b, ). It is anticipated that the integration of AI with VR and AR technologies will magnify their impacts within the metaverse (WEF, 2023b, ). These technologies facilitate the creation of hyper-realistic simulations, the development of digital entities and scenes, and enable deeper, more emotionally resonant interactions between users or with digital entities.
@@
Consequently, this technological expansion broadens the scope of threats like deepfakes (Tariq et al.,, 2023; Wu et al.,, 2023). In the absence of stringent regulatory oversight within the metaverse, users might exploit personalized identity markers – such as those derived from actual facial features and body forms – to commit identity theft or create synthetic identities. This poses significant risks of severe security and privacy violations, potentially leading to more profound forms of deception, defamation, and threats within the metaverse. For instance, the application of deepfake technology in virtual workplaces to fabricate digital clones of supervisors or colleagues could undermine internal trust, induce confusion by issuing deceptive work directives, or facilitate the dissemination of misinformation or propaganda, thereby impacting critical decisions (Tariq et al.,, 2023). Regrettably, existing deepfake detection methodologies are primarily tailored to address content in the physical world, such as images and videos, and do not yet adequately account for the unique challenges presented by the metaverse (Tariq et al.,, 2023). According to  Wu et al., (2023), there is a pressing need for deepfake recognition and prevention mechanisms specifically designed for the metaverse. Looking forward, addressing these challenges will necessitate a dual approach: developing robust technological solutions for identity verification and content authenticity, and establishing clear regulatory frameworks that delineate legal boundaries for virtual avatars and digital content, while addressing critical issues such as privacy protection and misuse prevention. Only through a coordinated effort can the metaverse be ensured to be a relatively safe and trustworthy virtual space.
@@ 4.5 Identity Fragmentation and Psychological Well-being Challenges
The principle of Expansiveness elucidates that the metaverse offers unparalleled opportunities for diverse and expansive identity representation, enabling individuals to transcend the limitations associated with the physical world. While this diversity in representation can enhance creativity and inclusivity, as exemplified by the Avatar Robot Cafe (Hatada et al.,, 2024) case where marginalized groups gained equal participatory opportunities, it also poses significant risks such as identity fragmentation and confusion. Recent research by  Yang et al., 2024b , involving a two-wave panel study on the VRChat platform, corroborates these concerns, indicating that while expansion of identity representation can boost self-esteem and life satisfaction, it may compromise these benefits when it leads to inconsistencies in self-concept. A poignant historical example is the 2000 incident where a child, after prolonged engagement with a video game, began to experience severe identity dissonance. Believing himself to be the game’s protagonist, he ventured out at night to “fight enemies and save the princess,” mimicking the game’s narrative. This blurring of reality and fantasy necessitated medical intervention, to help reorient his sense of identity and reality (Banerji,, 2023).
@@
As the metaverse progresses, addressing these psychological impacts and implementing protective measures for identity integrity, particularly among children and vulnerable populations, becomes imperative. Identity fragmentation within the metaverse can significantly impair users’ mental health and social relationships. The ability to easily create and switch between diverse, unintegrated identities may reduce reliance on real-world personas, exacerbating identity fragmentation issues (Yang et al., 2024b, ). These ‘consequence-free’ virtual spaces (Turkle,, 1999), which may diverge sharply from reality, pose specific risks. Such environments can lead to self-misalignment, where users perceive their virtual identities as superior, increasing their engagement with virtual communities and causing alienation from the physical world (Banerji,, 2023). This alienation can compel individuals, particularly children and adolescents, to increasingly immerse themselves in the metaverse, potentially leading to social withdrawal, depression, and antisocial behavior (Koohsari et al.,, 2023). Considering the known adverse effects of social media, video games, and mobile devices, the immersive and escapist nature of the metaverse requires scrutiny to prevent its potentially detrimental impacts on psychological and social well-being (Banerji,, 2023; Koohsari et al.,, 2023).
@@
Research on multiple cultural identities has shown that the mere presence of multiple identity representations is not inherently problematic; rather, the critical factor is relationships, integration, and management of these identity representations (Yang et al., 2024b, ). As the metaverse evolves and identity-related issues become more pronounced, further research and practical interventions are essential. Future studies should focus on assisting users in managing their multiple identity representations and preventing identity fragmentation to mitigate associated psychological and social challenges. Given that these concepts are still in their nascent stages, proactively recognizing and preparing for identity fragmentation is crucial for effectively addressing the emerging challenges of the metaverse.
@@ 5 Conclusion and Limitation
The metaverse, an emerging realm that merges virtual and real elements, significantly influences identity expression, social interaction, and human development trajectories. Central to the metaverse is the construction of identity, which involves the representation of personal traits, management of data footprint and social roles, identification of avatars and other digital entities, and economic elements. These aspects are intimately connected to users’ sense of belonging, privacy, security, and trust. Despite its significance, scholarly research on metaverse identities is still nascent and lacks a comprehensive theoretical framework. This paper conducts a critical review of the existing literature to explore the Metaverse Identity concept from both multidisciplinary and perspective viewpoints, identifying the research gap. Unlike traditional internet settings, the metaverse offers more immersive and varied life-like experiences, transforming identity from a mere theoretical construct into a tangible, experiential reality. This shift accentuates the tension between self-expression and privacy protection.
@@
Metaverse identities are characterized by their multilayered, dynamically evolving, and diverse nature, which grants users unprecedented freedom in self-expression but also poses substantial privacy and security risks. Addressing these complexities, this paper proposes two core principles: Equivalence and Alignment and Fusion and Expansiveness. The first principle argues for consistency between the metaverse and real-world identities in terms of behavioral norms and social standards, which is essential for developing conduct guidelines and protecting rights. The second principle stresses the need for deep integration and extensive expansion of metaverse identities, breaking beyond real-world constraints to meet diverse needs and promote inclusive participation. Effective governance in the metaverse necessitates a dynamic balance between these principles, ensuring fairness while encouraging diverse expressions and innovative developments. Our analysis identifies five key challenges to the development of metaverse identities: interoperability, legal boundaries, privacy and identity management issues, the risks associated with deep fakes & synthetic identities, and identity fragmentation impacting psychological health. Building on the principles outlined, this paper offers strategic recommendations to address these challenges.
@@
However, this study exhibits several significant limitations. Firstly, the representativeness of the case studies needs enhancement. This paper discusses instances such as virtual sexual assault and Avatar Robot Cafe, yet the case selection is narrow and largely confined to specific scenarios. Given the metaverse’s diverse and complex ecosystem, our analysis might not fully capture the spectrum of identity practices. Future research should broaden the scope of case studies to improve the findings’ generalizability and relevance. Secondly, the empirical foundation of this study requires strengthening. Metaverse identity research, still in its infancy, largely lacks comprehensive data on user behavior and psychological responses. Currently supported primarily by literature review and qualitative analysis, this study highlights the need for more rigorous empirical research. Future efforts should include observational, survey, and experimental methodologies to explore identity construction behaviors and perceptions in the metaverse more deeply, thereby solidifying the theoretical framework. Lastly, the theoretical principles introduced—Equivalence and Alignment, and Fusion and Expansiveness—need further refinement and operationalization. While these principles offer valuable insights for regulating identity in the metaverse, they are proposed from a broad perspective and remain abstract. The practical application and policy-making processes might encounter challenges in operationalizing these principles. Future research should refine these two principles and develop more detailed, actionable guidelines to inform the governance and design of the metaverse identity framework.
@@
Looking forward, the metaverse, as an emerging socio-technical ecosystem in the digital age (Lee et al.,, 2021; Wang et al.,, 2022; Lee et al., 2023b, ), is poised to significantly transform human life and developmental practices. The manner in which individuals construct and manage their identities between the physical world and the metaverse is critical for sustainable human development. Addressing this issue requires a unified approach from academia, industry, and policy sectors to promote sustained interdisciplinary and multi-perspective research, aimed at fostering a robust and inclusive metaverse environment. Such collaborative initiatives are essential to provide both intellectual and practical contributions to the evolution of this new era.